Can someone please work this out set by step please 2/3 x=1/3x+2
Answer by edjones(8007)

2/3 x=1/3x+2
These are all pretty simple if you eliminate the fraction by multiplying each side by the LCD.
2x=x+6
x=6
.
Ed